OpenStudy (anonymous):

A jewelry stand sells 5 different pairs of earrings. The earrings sell for $12, $9, $6, $8, and $10. What is the mean absolute deviation of the prices? Round to the nearest cent, if necessary.

OpenStudy (anonymous):

@Michele_Laino @dan815

OpenStudy (anonymous):

@confluxepic @Conqueror

OpenStudy (anonymous):

@confluxepic HELP

OpenStudy (michele_laino):

step#1 we have to find the mean value, of this set of values: $12, $9, $6, $8, and $10.

OpenStudy (michele_laino):

that mean value m is: \[m = \frac{{12 + 9 + 6 + 8 + 10}}{5} = ...?\]

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

9

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, you have to compute these quantities: \[\begin{gathered} 12 - 9 = ...? \hfill \\ 9 - 9 = ...? \hfill \\ 9 - 6 = ...? \hfill \\ 9 - 8 = ...? \hfill \\ 10 - 9 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

1.3 2.0 3.3 4.1 5.1

OpenStudy (michele_laino):

sorry, 12-9=3

OpenStudy (michele_laino):

please, continue

OpenStudy (anonymous):

umm i don't know what to do next :c

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have: \[\begin{gathered} 12 - 9 = 3 \hfill \\ 9 - 9 = 0 \hfill \\ 9 - 6 = 3 \hfill \\ 9 - 8 = 1 \hfill \\ 10 - 9 = 1 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

Those quantities, are the absolute deviations. now, your answer, namely the MAD, or mean absolute deviation is given by the subsequent computation: \[\large \frac{{3 + 0 + 3 + 1 + 1}}{5} = ...?\]

OpenStudy (anonymous):

8/5

OpenStudy (michele_laino):

and 8:5=1.6

OpenStudy (michele_laino):

so, your answer is 1.6

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

i have 4 more then i am done can you help please

OpenStudy (michele_laino):

ok!